OpenStudy (anonymous):

A=1/2bh solve for b Can someone show me the steps to solve this?

OpenStudy (lalaly):

2A=bh b=2A/h

OpenStudy (anonymous):

it means that b =2A/h

OpenStudy (anonymous):

a/h=1/2b (a/h)(2)=b b=(2a)/h

OpenStudy (anonymous):

so divide by h? i still don't quite get it.

OpenStudy (lalaly):

first u multiply both sides by 2 then you divide both sides by h

OpenStudy (anonymous):

she is ryt

OpenStudy (lalaly):

\[2 \times A=\frac{1}{\cancel{2}} bh \times \cancel{2}\]

OpenStudy (anonymous):

yea, you can do it in any order. You can divide by h first or divide by 1/2 first. Either way works because of the commutative property? Not sure if thats the right property but you can multiply or divide in any order.

OpenStudy (lalaly):

then divide by h

OpenStudy (anonymous):

ok thx for all of your answers :0 I am soooo confused with algebra!

OpenStudy (lalaly):

Its ok you just need practice, try solving more problems and if u ever get stuck post it here and me or anyone else can help you :D

OpenStudy (anonymous):

great, cause i have some more ??

OpenStudy (lalaly):

okay but first you gotta try them on ur own...

OpenStudy (anonymous):

Yea you'll get much more out of it if you think about them and work through them on your own first

OpenStudy (anonymous):

understand, i am old and never took algebra in h.s. spend many hours working on it!!!

OpenStudy (anonymous):

one thing meand - the way you have written the question is ambiguous - it can be read as ' a half times times bh' or 'one divided by 2bh'. to make sure theres no confusion if its the first one it is best written as (1/2)bh